Type or1. Suppose the Demand for down pillows is given by Qd = 200 - P and the supply is given by Qs = -10 + 6P a. Solve for the equ paste question here
Answer #1

1> Par eqaildie = 200-p 200-30 Srice 22 Pale elahi y a derad PED) in TPE0 d(o Q+ p$30 PED= d0-P) ap 30 2 00-P 130 2ou-30-30 (to 0.17 PES = dCas) ap Os -dc6p1o ap 6x30-1o 6 30 N. Snue Prce elwry op Sappiy (Pe) So male is greadas elastie as mprd